# Human-Machine Interfaces for Subsea Telerobotics: From Soda-straw to Natural Language Interactions

### v1

@@ Abstract
For more complex scenario, learning from/by demonstration (LfD/LbD) represents another collaborative paradigm, where the robot learns a knowledge base from operator-provided demonstrations and uses it to autonomously perform analogous tasks in the future. Unlike reward-driven reinforcement learning (RL), LfD relies on expert demonstration, resulting in behavior that closely mirrors the operator’s approach [179]. LfD has been extensively applied in subsea manipulation tasks, particularly for drilling and cutting [180], valve control [181], hot-stabbing [182], etc. A major limitation of many LfD approaches is they require optimal demonstration to learn, which is not always feasible, particularly for path planning. This is addressed by accounting for the variability in demonstration quality and combining Bayesian learning with probabilistic path planning [165]. Another challenge of LfD is its struggle to resolve discrepancies between the demonstration space and the robot’s operational space. To address this, Havoutis et al. [163] propose a probabilistic framework utilizing a hidden semi-Markov model to construct robust task representations. Their approach achieves a ten-fold reduction in manipulation task completion time using a simulated DexROV; however, they did not consider vehicle dynamics or external disturbances encountered in real-world environments. Among real-world projects, PANDORA [164] demonstrates the role of LfD in automating inspection and intervention tasks at subsea oil and gas facilities. The authors emphasize the necessity of a cognitive knowledge base enriched with semantic information about the workspace, as opposed to relying solely on pre-programmed trajectories or task-specific descriptions.
@@
In multi-robot systems, a directly teleoperated robot, known as the leader, works alongside one or more follower robots [173]. These followers possess some local autonomy and operate as a coordinated group. The human operator, controlling the leader, oversees the collective movement of the fleet [183]. Two fundamental challenges in such fleets are formation control and trajectory control [184]. The operator plays a crucial role in modulating the fleet’s velocity through the leader to ensure the followers maintain formation, avoid collisions, and stay synchronized with the leader. Such robotic fleets, also known as swarms, are useful for tasks requiring extended area coverage, dense scans, and rapid completion. Notable applications include search and rescue [161], structure inspection [6, 185], ship hull inspection [186, 187], large area mapping [9, 188, 189], cooperative localization [190], and continuous surveillance of underwater facilities [162]. The complexity of these tasks exceeds the capabilities of a single vehicle or imposes an excessive mental load on the operator. In contrast, a leader-follower multi-vehicle configuration can accelerate task completion, enhance vehicle cognition, and reduce the operator’s load.
@@
A recent survey by Zhou et al. [191] characterizes operators as independent from the swarm, limiting their role to mission initialization and maintenance; however, this perspective does not cover the full spectrum of multi-robot literature. For instance, several studies [192, 193] explore scenarios where follower robots autonomously estimate their relative pose while the operator directly controls the leader. Moreover, contemporary research emphasizes heterogeneous robot groups, unlike traditional homogeneous swarms, where all robots had the same characteristics and role [194]. In heterogeneous settings, the operator allocates tasks based on each robot’s specific capabilities and real-time operational status [195]. For marine applications, heterogeneous groups typically include an unmanned surface vehicle (USV) that acts as a communication relay between ROVs and the operator; an example is demonstrated in [6], consisting of one ASV and two SWIM-R ROVs for shallow water pipeline inspection. These configurations enhance mission reliability by leveraging the USV as an additional wireless communication channel.
14 4.3   Autonomous Systems
@@
As opposed to direct teleoperation and shared autonomy schemes discussed above, autonomous systems or AUVs operate independently, requiring no control commands from the operator. Subsequently, HMI for an AUV is primarily designed for status updates and mission management rather than direct control. However, even fully autonomous systems rely on human operators for mission planning, progress monitoring, and emergency intervention. This highlights the critical role of AUV interface design in ensuring the safety and effectiveness of autonomous missions.
@@
Generally, the interface includes a software console, a surface control module (SCM), and a transducer to maintain real-time communication with the AUV. Available real-time commands at the operator end are typically limited to starting or aborting the mission, requesting status updates, or intervening during critical moments. Contemporary works have integrated more advanced automated features into these interfaces. For instance, the MARES AUV interface automatically performs pre-mission procedures such as estimating mission duration and path length, testing communications, etc. [196]. Kaeli et al. [159] develop an anomaly detection framework that enables AUVs to focus on high-resolution scanning and transmit only the anomalous data to the operator for inspection, reducing data congestion and improving operational efficiency. A similar approach is reported for crack detection in subsea pipelines using a camera, laser, and multi-beam echosounder (MBES); the AUV detects potential cracks and sends the scanned images to an operator for further review [5]. The operator then requests close-up detailed examination if necessary, or continues the mission.
@@
Commercial AUV/UUV Platforms. Currently, the top application segments for AUV/UUVs are military and defense, offshore oil and gas drilling, oceanography, archaeology, and alternative energy harvesting [150]. Their operational costs are high due to factors including maintenance, repair, data acquisition, personnel training, and mission-specific logistics. Consequently, access to AUVs has been restricted to governments and their agencies, industry leaders, and a handful of academic and research labs. Examples of SOTA AUVs are L3Harris IVER (US), A9-E ECA (France), Teledyne Gavia (US), Kongsberg Hugin (Germany), SAAB Seaeye Sabertooth (Sweden), Exail A18 (France), and Atlas Elektronik SeaCat (Germany). Another lightweight low-cost category known as micro-AUV is gaining popularity for academic and research purposes; examples include NemoSens (France), Seaber Yuco series (France), and EcoSUB (UK). A detailed review of current AUV technologies, their challenges, and applications is outlined in [197].
15 5  Engineering Integration of Interactive Control
@@
16 5.1  Visual Feedback Technique
@@
17 5.2   Haptic Feedback Technique
@@
18 5.3   Pseudo-Haptic Feedback
@@
19 5.4   Other Feedback: Logs, Maps, Auditory Cues
@@
21 6 Challenges and Open Problems
The subsea telerobotics domain presents unique challenges due to the hostile and dynamic nature of the underwater environment. Key technical hurdles include limited FOV, latency in long-distance operations, hydrodynamic variations, uncertainty in state estimation, and the presence of dynamic obstacles within confined spaces. Open problems in this sector span the fields of robotics, low-light perception, AI, computer vision, human cognition, and ocean engineering. Fig. 12 highlights the promising directions for future research. Few prominent research areas and their open problems are discussed below.
@@
22 6.1   Human-Machine Dialogue in Natural Language
As natural language has been proven to be one of the most effortless control methods for teleoperation, the future goal is to enable real-time natural conversations, where the robot would possess a cognitive understanding of the goal and actively engage in decision-making. LLMs, particularly GPT models [237, 238], have successfully conducted real-time human-machine dialogues for various HRI settings [239, 240]. However, marine robots face the additional challenges of integrating sensory data and mission objectives into their decision-making process while executing a teleoperator’s command. ChatSim [121] and OceanGPT [122] have shown some progress in this regard, demonstrating limited dialogue capabilities and gathering vast knowledge of the ocean environment into the model, respectively. Further research can enable the robot to understand various instructions, reason logically, and propose alternative strategies using a cognitive understanding of a mission.
While language models can parse instructions and reason from a long conversation, they struggle with linguistic biases [125] such as variations in dialect/pronunciation, indirect commands, as well as mission-specific terminologies. Therefore, another research direction would be a risk assessment for voice-operated subsea missions and developing fail-safe mechanisms for the robot to identify, query, and correct ambiguous or potentially dangerous commands.
23 6.2   Optimizing Learning from Demonstration (LfD)
@@
In general, any reinforcement learning (RL) pipelines struggle with dynamic changes such as robot’s component failure or unseen terrain [242]. Extensive training (i.e., demonstrations) in diverse scenarios is a straightforward solution. However, crowdsourcing demonstrations are extremely challenging in harsh subsea environments, unlike ground/aerial scenarios [241]. This motivates multiple research directions. First, the robot should learn how to adapt to an unseen task using prior experience. Meta RL [243] and online RL [242] have shown progress in such model adaptation through reward-driven optimization. Researchers have utilized these approaches for improving subsea locomotion [244], path planning [245], and target tracking [246]. Second, realistic simulators featuring various underwater scenes can be a potential alternative for recording numerous task demonstrations. Third, algorithmic studies are needed to determine the minimum number of demonstrations required to achieve a desired performance threshold. Additionally, hybrid approaches that combine LfD and RL should be explored, where operator interventions during autonomous missions are treated as demonstrations or corrective feedback. Such interventions could potentially teach the robot to generalize knowledge, refine policies, and avoid similar errors in the future.
24 6.3   Hydrodynamic Sensing & Haptic Feedback Modeling
@@
Modeling haptic sensations from remote underwater settings presents multi-faceted challenges. First, the extreme pressure and complex hydrostatic/hydrodynamic forces limit the sensing accuracy [13]. Second, existing fluid dynamic models cannot accurately represent detailed information for different spatiotemporal scales of turbulent water flows. Third, translating high-fidelity hydrodynamic data at both the micro and macro levels in intuitive haptic feedback that effectively engages teleoperators’ sensory perception while maintaining ergonomics – remains as open problem.
@@
In recent years, researchers have explored materials such as Polycrystalline Nickel Titanium [247], Carbon Nanotubes [248], and Tungsten Disulfide Nanosheets [249], which exhibit flexible, hydrophobic, and corrosion-resistant properties. Doppler current profiler and artificial lateral line sensors have shown promising results for far-field and near-field flow measurements, respectively [99]. Other techniques such as Smoothed Particle Hydrodynamics (SPH) [250] and Position Based Dynamics (PBD) [251] are being explored for modeling free surface flows. Potential research directions in this regard are physics-driven learning [252] and generative models [253, 254] which are promising in turbulent flow modeling.
Existing vibrotactile sensation devices such as body suits provide multi-point vibrations but fall short in conveying the sense of fluid flow [33], which is essential for a fully immersive experience. As discussed earlier in Sec. 5, other kinesthetic feedback devices (e.g., exoskeleton) are expensive, less comfortable, and cannot mimic the unique underwater sensations. Therefore, engineering innovations for intuitive haptic sensation is another open problem in this domain.
25 6.4   Latency Mitigation for Long Distance Operation
@@
Mitigating the inherent latency in robot teleoperation has been a focal pursuit in contemporary research, given the pivotal implications these delays impose on overall system performance and operator experience. Researchers have tested USV-ROV pairs with LAN [255] and satellite connection [109] to enable faster communication. For existing tethered ROV systems, video prediction is an exciting research area. While it has gained some attention among the computer vision community [256], it has not been evaluated thoroughly for subsea telerobotics. More specifically, frame interpolation and future frame prediction can be a promising solution for the short-term disruption of video feed.
@@
On the other hand, optical-acoustic dual channels have established wireless communication up to a few hundred meters of depth [257]. RF signals have also been tested for wireless ROV control [258]; however, their application has been limited to shallow water applications. Unlike optical signals, RF waves are undisturbed by water turbulence, turbidity, and solar noise. Therefore, extending the working range of RF signals in a water medium would significantly improve operator-ROV communication in harsh environments. Compact magnetoelectric (ME) antennas [157] are also being explored as a potential modality for robot-robot and robot-teleoperator communications with promising early results.
@@
To account for long-term communication loss, predictive control algorithms and intent-aware robotic systems can offer potential solutions [259]. Predictive control algorithms can utilize historical data and real-time inputs to anticipate future states, while intent-aware systems would enable the robot to switch to autonomous mode and continue performing its tasks when a communication breakdown is imminent. Another approach is to reduce continuous manual input via intelligent delegation of control tasks to the autonomous subsystems within robots and mitigate the impacts of delays on the operator via sensory manipulation [260]. Although this technique has been widely exercised in rehabilitation literature [261], its potential in telerobotics remains under-explored. Further efforts are required from a broader research community that will combine supervisory controls, adaptive algorithms, and advanced sensory feedback into a cohesive framework for addressing the diverse challenges posed by teleoperation delays [260].
26 6.5   Real-World to/from Simulation
@@
Existing virtual simulators are capable of rendering simplified models of underwater workspace, including surface waves [130], chemical profiles [143], coral reefs [142], subsea rigs [134], and some dynamic obstacles [129]. However, they fail to represent the turbulent oceanic conditions and waterbody characteristics [262] in general. As compared earlier in Table 3, only a few simulators (e.g., UWSim and UUV Simulator) provide an extensive choice of sensors and diverse scenarios, making them suitable for testing navigation, SLAM, and path planning algorithms. Many platforms cannot simulate interactions with dynamic marine objects, which is critical for these tasks. The utility comparison of various simulators presented by Ciuccoli et al. [263] also highlights these limitations. Consequently, expanding the sensing capabilities and scenario diversity in simulators remains a pressing area for development.
Another issue is that most existing simulators cannot incorporate real-world data, leaving a gap between simulated scenarios and real-world missions. To address this, comprehensive real2sim frameworks should be developed, capable of translating real-time and historical sensor data from physical deployments into virtual platforms. For instance, project RUMI [28] leverages real-time video feeds to create an interactive terrain within a geo-referenced digital environment. Studies have demonstrated that such frameworks can enhance manipulation and navigation performance through improved realism [264]. Additionally, this approach narrows the gap between virtual and physical models, facilitating sim2real transfer of new algorithms and robotic systems developed in simulated scenarios [265].
Moreover, many existing DT models do not accurately represent their real-world pairs [263]. While they mimic the kinematic properties, onboard sensors and physical properties e.g., buoyancy and center of gravity are not replicated accurately. To improve realism, these design gaps in twin modeling need to be addressed and their interaction with dynamic environments should be extensively compared with real-world cases.
27 6.6   Adaptive Autonomy
@@
28 6.7   Brain-Machine Interfaces (BMI)
@@
29 7  Conclusions
The growing reliance on ROVs and AUVs for remote subsea expeditions has elevated the importance of designing advanced HMIs to ensure precise control and reliable interaction. Researchers have proposed various interfaces, from early vision-based consoles to more sophisticated, multimodal interfaces that integrate haptics, XR, and NL interactions – to enhance operator efficiency and comfort while equipping robots with greater autonomy and intelligence. In this paper, we reviewed the expansive body of literature on subsea HMI technologies, categorizing and analyzing their strengths, limitations, and applications. We identified unique challenges of subsea teleoperation and SOTA approaches to address them. The role of virtual interfaces and digital twins in lowering development costs, and facilitating operator training is also explored. We further examined SOTA shared autonomy frameworks and their role in bridging operators’ awareness with intelligent robots. Finally, we outlined several promising avenues for future research, including learning from demonstration, human-robot language interaction, cognitively aware systems, improved virtual interfaces, and beyond. Interdisciplinary collaboration across robotics, AI, and human cognitive science is required to meet the unique demands of the underwater domain and further advance HMI systems in subsea telerobotics.

### v2

@@ Abstract
For more complex scenario, learning from/by demonstration (LfD/LbD) represents another collaborative paradigm, where the robot learns a knowledge base from operator-provided demonstrations and uses it to autonomously perform analogous tasks in the future. Unlike reward-driven reinforcement learning (RL), LfD relies on expert demonstration, resulting in behavior that closely mirrors the operator’s approach [142]. LfD has been extensively applied in subsea manipulation tasks, particularly for drilling and cutting [143], valve control [144], hot-stabbing [145], etc. A major limitation of many LfD approaches is they require optimal demonstration to learn, which is not always feasible, particularly for path planning. This is addressed by accounting for the variability in demonstration quality and combining Bayesian learning with probabilistic path planning [134]. Another challenge of LfD is its struggle to resolve discrepancies between the demonstration space and the robot’s operational space. To address this, Havoutis et al. [132] propose a probabilistic framework utilizing a hidden semi-Markov model to construct robust task representations. Their approach achieves a ten-fold reduction in manipulation task completion time using a simulated DexROV; however, they did not consider vehicle dynamics or external disturbances encountered in real-world environments. Among real-world projects, PANDORA [133] demonstrates the role of LfD in automating intervention tasks at subsea oil and gas facilities. The authors emphasize the necessity of a cognitive knowledge base enriched with semantic information about the workspace, as opposed to relying solely on pre-programmed trajectories or task-specific descriptions.
@@
In multi-robot systems, a directly teleoperated robot, known as the leader, works alongside one or more follower robots [137]. These followers possess some local autonomy and operate as a coordinated group. The human operator, controlling the leader, oversees the collective movement of the fleet. Two fundamental challenges in such fleets are formation and trajectory control [146]. The operator plays a crucial role in modulating the fleet’s velocity through the leader to ensure the followers maintain formation, avoid collisions, and stay synchronized with the leader. Such robotic fleets, also known as swarms, are useful for tasks requiring extended area coverage and rapid completion. Notable applications include search and rescue [131], structure inspection [6], ship hull inspection [147], large area mapping [9], and cooperative localization [148]. The complexity of these tasks exceeds the capabilities of a single vehicle or imposes an excessive mental load on the operator. In contrast, a leader-follower multi-vehicle configuration can accelerate task completion, enhance vehicle cognition, and reduce the operator’s load.
@@
A recent survey by Zhou et al. [149] characterizes operators as independent from the swarm, limiting their role to mission initialization and maintenance; however, this perspective does not cover the full spectrum of multi-robot literature. Several studies [150, 151] present scenarios where follower robots autonomously estimate their relative pose while the operator directly controls the leader. Moreover, contemporary research emphasizes heterogeneous robot groups, unlike traditional homogeneous swarms, where all robots had the same characteristics and role [152]. In heterogeneous settings, the operator allocates tasks based on each robot’s specific capabilities and operational status [153]. Heterogeneous groups typically include an unmanned surface vehicle (USV) that acts as a communication relay between ROVs and the operator; an example is demonstrated in [6], consisting of one ASV and two SWIM-R ROVs for shallow water pipeline inspection. These configurations enhance mission reliability by leveraging the USV as an additional communication channel.
5.3 5.3   Autonomous Systems
@@
As opposed to direct teleoperation and shared autonomy schemes, autonomous systems or AUVs operate independently, requiring no control commands from the operator. Subsequently, HMI for an AUV is primarily designed for status updates and mission management rather than direct control. However, even fully autonomous systems rely on human operators for mission planning, monitoring, and intervention.
@@
Generally, the interface includes a software console, a surface control module (SCM), and a transducer to maintain real-time communication with the AUV. Available real-time commands at the operator end are typically limited to starting or aborting the mission, requesting status updates, or intervening during critical moments. Contemporary works have integrated more advanced automated features into these interfaces. For instance, the MARES AUV interface automatically performs pre-mission procedures such as estimating mission duration and path length, testing communications, etc. [154]. Kaeli et al. [129] develop an anomaly detection framework that enables AUVs to focus on high-resolution scanning and transmit only the anomalous data to the operator, reducing data congestion and improving efficiency. A similar approach is reported for crack detection in subsea pipelines using a camera, laser, and multi-beam echosounder; the AUV detects potential cracks and sends the scanned images to an operator for further review [5]. The operator then requests close-up detailed examination if necessary, or continues the mission.
@@
Current status & limitations. Currently, the top application segments for AUV/UUVs are military and defense, offshore oil and gas drilling, oceanography, archaeology, and alternative energy harvesting [155]. Their operational costs are high due to factors including maintenance, repair, data acquisition, personnel training, and mission-specific logistics. Consequently, access to AUVs has been restricted to governments and their agencies, industry leaders, and a handful of academic and research labs. Notable platforms such as Kongsberg Hugin (Germany), SAAB Seaeye Sabertooth (Sweden), and Exail A18 (France) have demonstrated long-range deployments for pre-scripted missions with limited surface intervention. Military-grade AUVs such as L3Harris IVER (US) and Atlas Elektronik SeaCat (Germany) offer surveying, intelligence gathering, mine disposal, etc. A lightweight, low-cost category known as micro-AUV is gaining popularity for academic and research purposes; examples include NemoSens (France), Seaber Yuco series (France), and EcoSUB (UK).
However, true autonomy – involving decision-making, adaptive planning, and high-level semantic understanding – remains an ongoing research challenge. We summarize the following limitations by reviewing the SOTA AUV literature.
  • Environmental uncertainty: Underwater environments are highly dynamic, with varying turbidity, current, and visibility, making perception and long-term localization unreliable.
  • Communication and energy constraints: Lack of real-time power and data transfer prevents operator supervision and long-term deployments.
  • Limited contextual reasoning: Current autonomous systems cannot reason about high-level goals or modify behavior based on task complexity or environmental cues.
  • Fault tolerance: Unlike teleoperated systems, AUVs lack immediate human oversight, which raises concerns for risk assessment and fault recovery.
Development trends. Recent academic and industrial trends in AUVs reflect a push toward scalability, adaptability, and improved intelligence. One notable direction is multi-AUV (swarm) systems for cooperative task execution (e.g., distributed mapping, search and rescue), which significantly enhance efficiency and spatial coverage [156]. Demand for low-cost and low-power AUVs is on the rise for academic research [157]. Additionally, learning-based techniques (e.g., LfD) are substituting explicit programming, as they allow AUVs to acquire complex skills by observing human operators [158]. Finally, the fusion of DT models with AUV platforms is gaining traction for risk mitigation, allowing for virtual planning and testing of mission strategies.
6 6. Engineering Integration of Interactive Control
@@
6.1 6.1  Visual Feedback Technique
@@
6.2 6.2   Haptic Feedback Technique
@@
6.3 6.3   Pseudo-Haptic Feedback
@@
6.4 6.4   Other Feedback: Logs, Maps, Auditory Cues
@@
7 7. Challenges and Open Problems
The subsea telerobotics domain presents unique challenges due to the hostile and dynamic nature of the underwater environment. Key technical hurdles include limited FOV, latency in long-distance operations, hydrodynamic variations, uncertainty in state estimation, and the presence of dynamic obstacles within confined spaces. Open problems in this sector span the fields of robotics, low-light perception, AI, computer vision, human cognition, and ocean engineering; see Fig. 12.
@@
7.1 7.1   Human-Machine Dialogue in Natural Language
As natural language has been proven to be one of the most effortless control methods for teleoperation, the future goal is to enable real-time natural conversations, where the robot would possess a cognitive understanding of the goal and actively engage in decision-making. LLMs, particularly GPT models [204], have successfully conducted real-time human-machine dialogues for various HRI settings [205, 206]. However, marine robots face the additional challenges of integrating sensory data into their decision-making process while executing an operator’s command. ChatSim [92] and OceanGPT [93] have shown some progress in this regard, demonstrating limited dialogue capabilities and gathering vast knowledge of the ocean environment into the model, respectively. Further research can enable robots to reason logically and propose alternative strategies.
While language models can parse instructions and reason from a long conversation, they struggle with linguistic biases [95] such as variations in dialect/pronunciation, indirect commands, as well as mission-specific terminologies. Therefore, another research direction would be a risk assessment for NL interfaces and developing error-handling mechanisms for the robot to identify, query, and correct ambiguous or potentially dangerous commands. In natural language generation (NLG) domain, errors are corrected via human feedback and reinforcement learning [207]. Lin et al. [208] compare three error handling strategies– reprompt, suggestion, and confirmation– across users of different age groups interacting with a chatbot. Among telerobotic NL interfaces, ChatSim [92] utilize closed perception-action loops to minimize error in GPT-generated mission plan. However, other interfaces are yet to adopt any structured error-handling mechanisms. To clarify ambiguous user queries, researchers use Similar Question Model (SQM) [209] that identifies ambiguous utterances by extracting features from past similar queries. Other approaches to reduce dialect and linguistic error include phonetic normalization, code mixing (i.e., mixing words from multiple language) [210] etc. To this end, these correction methods have been applied to search agents and AI assistants [211], hence their utility for robotic interfaces remains underexplored.
7.2 7.2   Optimizing Learning from Demonstration (LfD)
@@
In general, reinforcement learning (RL) pipelines struggle with dynamic changes such as robot’s component failure or unseen terrain [213]. Extensive training (i.e., demonstrations) in diverse scenarios is a straightforward solution. However, crowdsourcing demonstrations are extremely challenging in harsh subsea environments, unlike ground/aerial scenarios [212]. This motivates multiple research directions. First, the robot should learn how to adapt to an unseen task using prior experience. Meta RL [214] and online RL [213] have shown progress in such model adaptation through reward-driven optimization. Researchers have utilized these approaches for improving subsea locomotion [215] and path planning [216]. Second, realistic simulators featuring various scenes can be a potential alternative for recording numerous demonstrations. Third, algorithmic studies can reveal the minimum number of demonstrations required to achieve a desired performance threshold. Additionally, hybrid approaches (LfD and RL) should be explored, where operator interventions during autonomous missions are treated as demonstrations or corrective feedback. Such interventions could potentially teach the robot to generalize knowledge and refine policies.
7.3 7.3   Hydrodynamic Sensing & Haptic Feedback Modeling
@@
Modeling haptic sensations from remote underwater settings presents multi-faceted challenges. First, the extreme pressure and complex hydrostatic/hydrodynamic forces limit the sensing accuracy [58]. Second, existing fluid dynamic models cannot accurately represent different spatiotemporal scales of turbulent water flows. Third, translating high-fidelity hydrodynamic data at both micro and macro levels in intuitive haptic feedback that engages operators’ sensory perception while maintaining ergonomics – remains an open problem.
@@
In recent years, researchers have explored new sensing materials such as Polycrystalline Nickel Titanium [217] and Tungsten Disulfide Nanosheets [218], which exhibit hydrophobic and corrosion-resistant properties. Doppler current profiler and artificial lateral line sensors have shown promising results for far-field and near-field flow measurements, respectively [69]. Other techniques such as Smoothed Particle Hydrodynamics (SPH) [219] and Position Based Dynamics (PBD) [220] are being explored for modeling free surface flows. Potential research directions in this regard are physics-driven learning [221] and generative models [222], which show promise in turbulent flow modeling.
7.4 7.4   Latency Mitigation for Long Distance Operation
@@
Mitigating the inherent latency in robot teleoperation has been a focal pursuit in contemporary research, given its impact on overall system performance and operator experience. Researchers have tested USV-ROV pairs with LAN [223] and satellite connection [79] for faster communication. For existing tethered ROV systems, video prediction and frame interpolation is an exciting research areas. While it has gained some attention among the computer vision community [224], it has not been evaluated thoroughly for telerobotics.
@@
On the other hand, optical-acoustic dual channels have demonstrated wireless communication up to a few hundred meters of depth [225]. RF signals have also been tested for wireless ROV control [226]; however, their application has been limited to shallow water applications. Unlike optical signals, RF waves are undisturbed by water turbulence, turbidity, and solar noise. Therefore, extending the working range of RF signals in a water medium would significantly improve operator-ROV communication in harsh environments. Compact magnetoelectric (ME) antennas [126] are also being explored as a potential modality for robot-robot and robot-teleoperator communications with promising early results.
@@
To account for long-term communication loss, predictive control algorithms and intent-aware robotic systems can offer potential solutions [227]. Predictive control utilizes historical data and real-time inputs to anticipate future states, while intent-aware systems enable the robot to switch to autonomous mode and continue performing its tasks when a communication breakdown is imminent. Another approach is to reduce continuous manual input via intelligent delegation of control tasks to mitigate the impacts of delays on the operator via sensory manipulation [228]. Although this technique has been widely exercised in rehabilitation literature [229], its potential in telerobotics remains under-explored. Further efforts are required to combine supervisory controls, adaptive algorithms, and advanced sensory feedback into a cohesive framework for addressing the challenges posed by teleoperation delays [228].
7.5 7.5   Real-World to/from Simulation
@@
Existing virtual simulators are capable of rendering simplified models of underwater workspace, including surface waves [100], chemical profiles [116], coral reefs [113], subsea rigs [101], and some dynamic obstacles [117]. However, they fail to represent the turbulent oceanic conditions and waterbody characteristics [230] in general. As compared earlier in Table 3, only a few simulators (e.g., UWSim and UUV Simulator) provide an extensive choice of sensors and diverse scenarios to support navigation, SLAM, and planning algorithms. Many platforms cannot simulate interactions with dynamic marine objects, which is critical for the aforementioned tasks. The utility comparison of various simulators [231] also highlights these limitations. Expanding the sensor suites and scenario diversity in simulators remains an open research area.
Another limitation of existing simulators is their inability to incorporate real-world data, leaving a gap between simulation and real-world missions. Robust real2sim frameworks are needed to translate real-time and historical sensor data into virtual environments. For instance, project RUMI [166] leverages real-time video feeds to create an interactive terrain within a digital environment to improve realism [232]. Eventually, this approach may facilitate sim2real transfer of algorithms developed in simulation [233]. Moreover, many digital shadow/twin models fail to accurately replicate real-world pairs, lacking fidelity in sensor modeling and environmental dynamics [231]. Addressing these gaps is essential for achieving operational realism and system performance of simulated environments.
7.6 7.6   Adaptive Autonomy
@@
7.7 7.7   Brain-Machine Interfaces (BMI)
@@
8 8. Conclusions
The increasing dependence on ROVs and AUVs for subsea operations underscores the critical need for advanced HMIs that enable precise control, robust interaction, and mission success. In this review, we synthesized a broad spectrum of literature on subsea HMI technologies, systematically categorizing and analyzing their capabilities, limitations, and application domains. We identified the unique operational challenges inherent to subsea teleoperation and examined state-of-the-art approaches designed to mitigate them. Special attention was given to the role of virtual interfaces and digital twins in reducing development costs and enhancing operator training. Furthermore, we explored shared autonomy frameworks that integrate operator intent with autonomous decision-making to improve situational awareness and system responsiveness. The review concludes by highlighting several promising directions for future research, including learning from demonstration, natural language-based interaction, and cognitively adaptive systems. Addressing these challenges will require sustained interdisciplinary collaboration across robotics, artificial intelligence, and cognitive science to advance the next generation of HMI systems for subsea telerobotics.